OpenStudy (k_lynn):

Medal and Fan! Portia drew a line on the coordinate plane that passed through the points (-3,19) and (2,9). Which equation represents the line that Portia drew? Please Explain

OpenStudy (k_lynn):

Here's the picture

OpenStudy (k_lynn):

@StudyGurl14

OpenStudy (studygurl14):

Use the slope formula. Do you know what that is?

OpenStudy (k_lynn):

This is slope

OpenStudy (k_lynn):

I know it's either c or d, but I don't know which one @StudyGurl14

OpenStudy (studygurl14):

yep, that's the formula... And both C and D are incorrect

OpenStudy (k_lynn):

then I have no Idea how to solve this.

OpenStudy (studygurl14):

\(\large m=\Large\frac{y_2-y_2}{x_2-x_1}\rightarrow\frac{9-19}{2-(-3)}\rightarrow\frac{-10}{5}\rightarrow ?\)